# A Survey of Efficient Reasoning for Large Reasoning Models: Language, Multimodality, and Beyond

### v1

@@ 4.3 Summary and Outlook
In this section, we have reviewed methods that enhance efficient reasoning in LLMs via SFT. Key approaches include reasoning chain compression, where models reduce the length and complexity of CoT sequences through techniques like token budget control Han et al. (2024), self-training Munkhbat et al. (2025), and dynamic token skipping Xia et al. (2025). These methods optimize reasoning without losing accuracy, especially for tasks with simpler or more parallelizable structures. Another major approach is latent space reasoning, where explicit CoT steps are replaced by continuous representations in hidden states. Techniques like Coconut Hao et al. (2024), CCoT Cheng and Van Durme (2024), and CODI Shen et al. (2025c) use latent representations to improve efficiency and reduce token overhead. Innovations such as Token Assorted Su et al. (2025), SoftCoT Xu et al. (2025b), and Heima Shen et al. (2025a) further enhance reasoning by mixing latent and text tokens, reducing memory usage and computational cost.
As LLMs evolve, future research on efficient reasoning may focus on refining latent-space methods and developing models that are both more flexible and scalable. A promising direction is the integration of explicit and implicit reasoning processes, allowing models to dynamically switch between different reasoning strategies based on task complexity Yang et al. (2024c). Additionally, exploring multi-modal latent reasoning, where models leverage both textual and visual data, could enhance their reasoning abilities Talmor et al. (2020); Sun et al. (2024b). Research into adaptive curriculum learning strategies will further improve model flexibility, enabling them to handle more complex reasoning tasks Kong et al. (2021). Finally, addressing CoT unfaithfulness and improving alignment between explicit reasoning paths and internal latent representations will be critical for ensuring model reliability across diverse tasks Li et al. (2024a); Arcuschin et al. (2025).
@@ 5 Efficient Reasoning with Reinforcement Learning
| Name | RL Algorithm | On/Off Policy | Reward Function |
| O1-Pruner Luo et al. (2025a) | PPO | Off-policy | |
[…]
DeepSeek-R1 has demonstrated that RL can effectively guide language models to develop deep reasoning abilities, marking a significant advancement in enhancing models’ cognitive capabilities. This breakthrough naturally suggests that RL frameworks could potentially be a more direct and intuitive approach to improving reasoning efficiency. Inspired by this potential, several research efforts have emerged to explore the intersection of reinforcement learning and reasoning efficiency, focusing on developing more sophisticated RL-based training strategies to reduce or control the token usage of reasoning trajectories while maintaining accuracy. Depending on the approach to modeling reasoning length, we classify current approaches that leverage RL for efficient reasoning into two distinct categories, as shown in Figure 7.
5.1 Efficient Reinforcement Learning with Length Reward
Introducing a length reward alongside the rule-based reward provides a natural approach to efficiency. As shown in Table 2, several research efforts have explored the effectiveness of this method. Some studies establish the generation token budget, by either formulating correlations between task difficulty and the generation length or specifying in prompt. DAST Shen et al. (2025b) introduces Difficulty-Adaptive Slow-Thinking that empowers models to modulate CoT length based on problem complexity autonomously. They first define a Token Length Budget (TLB) metric, quantifying the difficulty of a problem by its success rate, then leverage length-aware reward shaping and length preference optimization to realize DAST. LCPO Aggarwal and Welleck (2025) controls the length budget by introducing a target length instruction in the prompt, i.e., “Think for tokens”, and designs a target-aware length reward that penalizes the length violation.
@@ 5.1 Efficient Reinforcement Learning with Length Reward
Other approaches incorporate the length reward normalized by a baseline budget. O1-Pruner Luo et al. (2025a) designs an efficient fine-tuning method that begins by estimating the LLM’s baseline performance through presampling from its reference model. Different from O1-Pruner, Arora and Zanette (2025) introduce a length penalty normalized in the per-prompt group. This strategy encourages the model to produce correct responses with a minimum amount of tokens while maintaining that correct responses are always preferred over incorrect ones. Kimi 1.5 technical report Team et al. (2025) discusses an observation of the overthinking phenomenon and introduces a length reward to restrain the rapid growth of token length. The length reward they defined is a normalized length factor compared to the maximum and minimum lengths of the different generated solutions.
Yeo et al. (2025) analyzes RL design choices for reasoning, revealing that extremely long CoT reasoning (approaching context limits) paradoxically reduces accuracy. Their proposed cosine reward function provides intuitive guidance - gradually increasing rewards for meaningful reasoning steps while penalizing excessive length. They also identify “length hacking”, where models artificially extend reasoning on difficult questions through repetition rather than genuine problem-solving, highlighting the challenge of aligning length-based rewards with actual reasoning quality.
5.2 Efficient Reinforcement Learning without Length Reward
There are also various approaches addressing this challenge without relying on explicit length rewards. These include innovative reformulations of the reinforcement learning framework, optimization methods that balance preferences between concise and detailed chain-of-thought solutions, and the identification of GRPO inductive bias.
MRT Qu et al. (2025b) formulates the test-time optimization problem into a meta-reinforcement learning problem. It divides the generation into episodes, requiring the model to estimate answers after each episode. This approach creates a natural trade-off: the model receives rewards for both correct early predictions (exploitation) and for continuing to refine its reasoning when uncertain (exploration), effectively learning to balance speed and accuracy across different compute budgets.
Methods that balance preferences include IBPO Yu et al. (2025b), which approaches budget awareness by framing it as utility maximization rather than directly controlling response length. Their key insight is categorizing responses into standard and extended reasoning groups, and optimizing the distribution between these groups. Chen et al. (2025c) employs heuristics such as First-Correct Solutions (FCS) and Greedy Diverse Solutions (GDS) to generate preference data for offline policy optimization using preference optimization methods (i.e., DPO, RPO, and SimPO) to control reasoning path lengths. Finally, Liu et al. (2025c) study the increasing length of reasoning trajectory from the inductive bias of RL algorithms. They claim that the token-level averaging of policy loss and standard deviation normalization in GRPO advantage introduces inductive bias to longer reasoning trajectories. To address this, they introduce Dr.GRPO to remove such bias and achieve more efficient RL scaling.
@@ 5.3 Summary and Outlook
The existing approaches commonly employ reinforcement learning techniques to optimize the trade-off between reasoning depth and token efficiency, with shared focuses on reward engineering that penalizes excessive length while preserving accuracy. Most methods formulate the challenge as a constrained optimization problem, designing specialized reward functions that balance accuracy with length penalties, though they differ in how they quantify and enforce these constraints, ranging from explicit budget targets to adaptive difficulty-based adjustments.
Despite significant progress, insights into why reinforcement learning inefficiently scales with sequence length are less explored Liu et al. (2025c), making principled solutions elusive. In addition, existing work focuses on verifiable tasks, such as reasoning and math. A promising direction is the efficiency of RL with general tasks using reward models or multi-modal tasks. Furthermore, as Qu et al. (2025b) initially explored, developing RL methods where more tokens consumption leads to better performance, represents an exciting direction. Current research primarily focuses on long CoT reasoning in o1/R1-like models, while efficient RL methods for alternative reasoning structures, such as parallel search, tree of thoughts, or graph of thoughts, remain largely unexplored.
@@ 6.1 Pretraining with Latent Space
Pretraining in latent space has recently garnered attention as a means of improving reasoning efficiency in LRMs. Instead of relying on traditional token-based approaches, these methods explore continuous representations, enhancing the depth of model understanding and efficiency Hao et al. (2024); Deng et al. (2024). Several approaches have emerged that differ in their treatment of the latent space and pretraining strategies (Sun et al., 2024d).
Byte Latent Transformer (BLT) processes raw bytes using dynamically sized patches rather than fixed tokens, thereby reducing computational overhead and improving robustness against noisy and multilingual inputs Pagnoni et al. (2024). By grouping bytes based on entropy rather than relying on predetermined tokenizations, BLT achieves better scalability and long-tail generalization Song et al. (2025a). Large Concept Models (LCMs) extend this paradigm at a higher semantic level. In LCMs, abstract concepts that often correspond to complete sentences or speech utterances serve as the primary processing units. They utilize pre-existing sentence embedding spaces SONAR Duquenne et al. (2023) and autoregressive sentence prediction to achieve modality- and language-agnostic performance The et al. (2024).
CoCoMix (Continuous Concept Mixing) integrates discrete token prediction with continuous concept vectors derived from sparse autoencoders (SAEs) Bricken et al. (2023); Cunningham et al. (2023). By interleaving these continuous representations into the hidden states during training, CoCoMix improves sample efficiency and enriches the model’s capacity for higher-order abstraction, which is beneficial for tasks such as summarization and logical reasoning Tack et al. (2025). Additionally, latent thought vectors (LTMs) were introduced to probabilistically guide token generation via cross-attention mechanisms Kong et al. (2025). LTMs are sampled from a latent prior and offer a flexible framework that allows adjustments of both inference steps and latent vector dimensions to optimize performance. This method improves sample efficiency and scalability compared to traditional autoregressive models.
Despite these advances, challenges remain. The reliance on implicit reasoning shortcuts may impede true stepwise reasoning, particularly in complex, multi-step tasks Lin et al. (2025). Future research may focus on refining these latent space pretraining methods to ensure consistency and accuracy across a broader range of reasoning tasks.
@@ 6.2 Subquadratic Attention
Among subquadratic attention mechanisms, linear sequence modeling techniques Sun et al. (2025c), such as linear attention, state space models (SSMs), and linear RNNs, emerge as effective alternatives to traditional self-attention. Sparse attention also presents a viable solution by selectively focusing on a subset of tokens, further improving computational efficiency. In the following sections, we will explore these subquadratic attention mechanisms in detail. Additionally, we will discuss recent linearization methods that transform pre-trained transformer-based model weights into linear recurrent structures, allowing for efficient inference while maintaining the knowledge embedded in the original transformer models.

### v2

@@ 4.3 Summary and Outlook
In this section, we have reviewed methods that enhance efficient reasoning in LLMs via SFT. Key approaches include reasoning chain compression, where models reduce the length and complexity of CoT sequences through techniques like token budget control han2024token, self-training munkhbat2025selftrainingelicitsconcisereasoning, and dynamic token skipping xia2025tokenskip. These methods optimize reasoning without losing accuracy, especially for tasks with simpler or more parallelizable structures. Another major approach is latent space reasoning, where explicit CoT steps are replaced by continuous representations in hidden states. Techniques like Coconut hao2024training, CCoT cheng2024compressed, and CODI shen2025codicompressingchainofthoughtcontinuous use latent representations to improve efficiency and reduce token overhead. Innovations such as Token Assorted su2025token, SoftCoT xu2025softcot, and Heima shen2025efficient further enhance reasoning by mixing latent and text tokens, reducing memory usage and computational cost.
As LLMs evolve, future research on efficient reasoning may focus on refining latent-space methods and developing models that are both more flexible and scalable. A promising direction is the integration of explicit and implicit reasoning processes, allowing models to dynamically switch between different reasoning strategies based on task complexity yang-etal-2024-large-language-models. Additionally, exploring multi-modal latent reasoning, where models leverage both textual and visual data, could enhance their reasoning abilities talmor2020leap; sun2024surf. Research into adaptive curriculum learning strategies will further improve model flexibility, enabling them to handle more complex reasoning tasks kong2021adaptive. Finally, addressing CoT unfaithfulness and improving alignment between explicit reasoning paths and internal latent representations will be critical for ensuring model reliability across diverse tasks li2024towards; arcuschin2025chain.
@@ 5 Efficient Reasoning with Reinforcement Learning
| Name | RL Alg. | Type | Reward Function |
| O1-Pruner luo2025o1 | PPO | Off | |
| Efficient arora2025traininglanguagemodelsreason | PPO | On | |
| L1 aggarwal2025l1controllinglongreasoning | GRPO | On | |
| Kimi-1.5 kimiteam2025kimik15scalingreinforcement | OPMD | On | |
| Demystifying yeo2025demystifyinglongchainofthoughtreasoning | PPO | On | |
| Redundancy Hong2025ReconsideringOP | PPO | On | |
| VSRM Yue2025PromotingER | PPO/Rein+ | On | |
| SmartThinker He2025SmartThinkerLT | SCPO | On | |
| Train Long | GRPO | On | |
| , Think Short Hammoud2025TrainLT | |||
| ShorterBetter yi2025shorterbetter | GRPO | On | |
| GFPO Shrivastava2025SampleMT | GRPO | On | |
| DAST shen2025dastdifficultyadaptiveslowthinkinglarge | SimPO | Off | |
| Aware First, | GRPO | On | |
| Think Less chen2025awarefirstthinkless | |||
| SABER zhao2025saber | GRPO | On | |
| ASRR zhang2025when | GRPO | On | |
| ARM wu2025arm | Ada-GRPO | On | |
| ACPO cheng2025incentivizingdualprocess | PPO/GRPO | On | |
| LASER liu2025learnreasonefficientlyadaptive | PPO | On | |
| ConciseRL dumitru2025conciserl | PPO | On | |
| ThinkPrune hou2025thinkprune | PPO | On | |
| TWYN yang2025thinkwhenyouneed | GRPO | On | |
| CANON chen2025conditionaladvantage | GRPO | On |
DeepSeek-R1 has demonstrated that reinforcement learning can effectively guide language models to develop strong reasoning capabilities, representing a notable advance in improving model cognition zhang2025survey. This progress naturally raises the question of whether reinforcement learning can also serve as a principled and intuitive framework for improving reasoning efficiency, rather than solely reasoning accuracy. Motivated by this possibility, a growing body of work has begun to explore the intersection of reinforcement learning and efficient reasoning, with the goal of reducing or better controlling token consumption along reasoning trajectories while maintaining task performance. Depending on how reasoning length is modeled and regulated during training, we categorize existing reinforcement learning approaches for efficient reasoning into two broad classes, as illustrated in Figure 7.
Introducing a length-aware reward alongside rule-based correctness signals provides a natural mechanism for improving reasoning efficiency within reinforcement learning. As summarized in Table 2, numerous studies have investigated how explicit control over generation length can mitigate overthinking while preserving accuracy. These methods differ in how length is quantified and enforced, ranging from fixed or difficulty-adaptive token budgets to finer-grained step-level and reasoning-mode–aware controls. To systematize this rapidly evolving literature, we organize reinforcement learning methods with explicit length rewards along several key design dimensions and discuss representative approaches under the following taxonomy.
5.1 Trajectory-level Budgeted Length Rewards
@@
yeo2025demystifyinglongchainofthoughtreasoning analyzes RL design choices for reasoning, revealing that extremely long CoT reasoning (approaching context limits) paradoxically reduces accuracy. Their proposed cosine reward function provides intuitive guidance—gradually increasing rewards for meaningful reasoning steps while penalizing excessive length. They also identify “length hacking”, where models artificially extend reasoning on difficult questions through repetition rather than genuine problem-solving, highlighting the challenge of aligning length-based rewards with actual reasoning quality. LCPO aggarwal2025l1controllinglongreasoning controls the length budget by introducing a target length instruction in the prompt, i.e., “Think for tokens”, and designs a target-aware length reward that penalizes the length violation. Other approaches incorporate the length reward normalized by a baseline budget. O1-Pruner luo2025o1 designs an efficient fine-tuning method that begins by estimating the LLM’s baseline performance through presampling from its reference model. Different from O1-Pruner, arora2025traininglanguagemodelsreason introduce a length penalty normalized in the per-prompt group. This strategy encourages the model to produce correct responses with a minimum amount of tokens while maintaining that correct responses are always preferred over incorrect ones. Kimi 1.5 technical report kimiteam2025kimik15scalingreinforcement discusses an observation of the overthinking phenomenon and introduces a length reward to restrain the rapid growth of token length. The length reward they defined is a normalized length factor compared to the maximum and minimum lengths of the different generated solutions.
These methods share a common formulation that treats reasoning length as a trajectory-level resource and frames efficient reasoning as a constrained optimization problem, where the objective is to remain within a soft or target budget while preserving correctness.
5.2 Redundancy-aware and Step-level Length Rewards
Motivated by the limitations of purely trajectory-level penalties, several works refine length rewards by introducing step-level or redundancy-aware signals. Researchers Hong2025ReconsideringOP decompose overthinking into internal redundancy within the first correct reasoning trace and external redundancy caused by unnecessary continuation after reaching a correct solution, and propose a dual-penalty reinforcement learning objective targeting both forms. VSRM Yue2025PromotingER introduces verifiable step-level reward signals that explicitly evaluate the utility of intermediate reasoning steps. SmartThinker He2025SmartThinkerLT further demonstrates that global length rewards may over-compress critical reasoning steps while preserving redundancy elsewhere, and proposes step-level length control based on reasoning-step importance.
5.3 Adaptive Budget Allocation
Beyond fixed or normalized length budgets, a line of work observes that the optimal amount of reasoning is inherently input-dependent, and explicitly trains models to allocate reasoning effort adaptively through length-aware reinforcement learning. Train Long, Think Short Hammoud2025TrainLT formulates efficient reasoning as a curriculum learning problem under GRPO, where training begins with relatively loose reasoning budgets and progressively tightens length constraints, encouraging models to maintain correctness under increasingly strict budget limits. ShorterBetter yi2025shorterbetter further leverages multi-sample training-time exploration by defining a shortest-correct target among sampled trajectories and training the model to prefer minimal-length correct reasoning. GFPO Shrivastava2025SampleMT improves test-time efficiency by trading additional training-time sampling for shorter inference trajectories, filtering candidate rollouts using reward-per-token and length-normalized efficiency signals during RL optimization. DAST shen2025dastdifficultyadaptiveslowthinkinglarge introduces Difficulty-Adaptive Slow-Thinking that empowers models to modulate CoT length based on problem complexity autonomously. They first define a Token Length Budget (TLB) metric, quantifying the difficulty of a problem by its success rate, then leverage length-aware reward shaping and length preference optimization to realize DAST. Aware First, Think Less chen2025awarefirstthinkless introduces a dynamic boundary self-awareness mechanism that enables models to identify sufficient reasoning states and terminate generation early, effectively allocating reasoning budgets based on internal confidence signals rather than static length constraints.
5.4 Reasoning Mode Switching
Complementary to continuous budget allocation, another line of work explicitly models discrete reasoning modes and trains models to switch among them under length-aware reinforcement learning objectives. SABER zhao2025saber introduces discrete reasoning budget tiers corresponding to different inference modes (e.g., no-thinking, fast-thinking, and deep-thinking), and trains models to switch among these modes using length-constrained reinforcement learning. ASRR zhang2025when proposes an accuracy-aware length regulation mechanism that dynamically allocates additional reasoning steps when needed, revealing an internal self-recovery phenomenon during long-chain reasoning. ARM wu2025arm further generalizes reasoning mode switching by treating different reasoning formats (e.g., direct answer, short CoT, long CoT, and code-based reasoning) as explicit actions, and applies reinforcement learning to select formats that balance correctness and reasoning cost. Incentivizing Dual Process Thinking cheng2025incentivizingdualprocess explicitly distinguishes fast (System 1) and slow (System 2) reasoning processes, and introduces process-aware incentives that encourage models to invoke long-form reasoning only when necessary.
5.5 Alternative Efficiency Signals
Beyond directly penalizing raw token counts, several approaches incorporate higher-level efficiency signals that still explicitly encode length or conciseness preferences within the reinforcement learning objective. LASER liu2025learnreasonefficientlyadaptive provides a unifying perspective on length-based reward shaping, and proposes step-function as well as difficulty-aware shaping strategies that adjust reward according to reasoning length and task complexity. ConciseRL dumitru2025conciserl replaces explicit length penalties with an LLM-judged conciseness score as the reinforcement signal, using semantic conciseness as a proxy for length efficiency. ThinkPrune hou2025thinkprune iteratively tightens hard token limits during reinforcement learning, gradually pruning unnecessary reasoning steps while preserving task performance. Think When You Need yang2025thinkwhenyouneed constructs comparative length–quality rewards that explicitly balance answer correctness against reasoning length, encouraging models to reason only when additional tokens improve outcomes. CANON chen2025conditionaladvantage further departs from explicit length modeling by conditioning advantage estimation on solution attributes, improving credit assignment in long reasoning trajectories and indirectly encouraging more concise and efficient reasoning.
5.6 Implicit Length Bias
In contrast to methods that explicitly impose length or budget constraints, another line of work improves reasoning efficiency without introducing direct length-related rewards. These approaches exploit implicit biases arising from reinforcement learning formulations, preference modeling, or training dynamics, allowing efficient reasoning behaviors to emerge naturally rather than being explicitly enforced.
MRT qu2025optimizingtesttimecomputemeta reformulates test-time computation as a meta-reinforcement learning problem. It decomposes generation into multiple episodes and requires the model to produce intermediate answers after each episode. This design induces an inherent trade-off between exploitation and exploration: the model is rewarded for making correct early predictions while also being incentivized to continue reasoning when uncertainty remains, thereby learning to allocate computation adaptively across different test-time budgets without explicit length penalties. A related direction focuses on balancing preferences between concise and extended reasoning trajectories. IBPO yu2025think frames budget awareness as a utility maximization problem rather than directly constraining response length. Specifically, it categorizes responses into standard and extended reasoning modes and optimizes the preference distribution between them. Similarly, chen2025think23overthinkingo1like employs heuristics such as First-Correct Solutions (FCS) and Greedy Diverse Solutions (GDS) to construct preference data for offline policy optimization, leveraging preference-based methods (e.g., DPO, RPO, and SimPO) to implicitly regulate reasoning path lengths. Beyond algorithmic design, recent work has begun to analyze the intrinsic inductive biases of reinforcement learning algorithms themselves. liu2025understanding show that the increasing length of reasoning trajectories can arise from token-level loss averaging and standard deviation normalization in GRPO advantage estimation, which together introduce a systematic bias toward longer generations. To mitigate this effect, they propose Dr.GRPO, which removes such bias and achieves more efficient scaling behavior without modifying the reward function.
@@ 5.7 Summary and Outlook
The existing approaches commonly employ reinforcement learning techniques to optimize the trade-off between reasoning depth and token efficiency, with a shared focus on reward engineering that penalizes excessive length while preserving accuracy. Most methods formulate the challenge as a constrained optimization problem, designing specialized reward functions that balance accuracy with length penalties, though they differ in how they quantify and enforce these constraints, ranging from explicit budget targets to adaptive difficulty-based adjustments.
Despite significant progress, insights into why reinforcement learning inefficiently scales with sequence length are less explored liu2025understanding, making principled solutions elusive. In addition, existing work focuses on verifiable tasks, such as reasoning and math. A promising direction is the efficiency of RL with general tasks using reward models or multi-modal tasks. Furthermore, as qu2025optimizingtesttimecomputemeta initially explored, developing RL methods where more tokens consumption leads to better performance, represents an exciting direction. Current research primarily focuses on long CoT reasoning in o1/R1-like models, while efficient RL methods for alternative reasoning structures, such as parallel search, tree of thoughts, or graph of thoughts, remain largely unexplored.
@@ 6.1 Pretraining with Latent Space
Pretraining in latent space has recently garnered attention as a means of improving reasoning efficiency in LRMs. Instead of relying on traditional token-based approaches, these methods explore continuous representations, enhancing the depth of model understanding and efficiency hao2024training; deng2024explicit. Several approaches have emerged that differ in their treatment of the latent space and pretraining strategies (sun2024co2).
Byte Latent Transformer (BLT) processes raw bytes using dynamically sized patches rather than fixed tokens, thereby reducing computational overhead and improving robustness against noisy and multilingual inputs pagnoni2024byte. By grouping bytes based on entropy rather than relying on predetermined tokenizations, BLT achieves better scalability and long-tail generalization song2025headtailbalancedrepresentation. Large Concept Models (LCMs) extend this paradigm at a higher semantic level. In LCMs, abstract concepts that often correspond to complete sentences or speech utterances serve as the primary processing units. They utilize pre-existing sentence embedding spaces SONAR duquenne2023sonar and autoregressive sentence prediction to achieve modality- and language-agnostic performance the2024large.
CoCoMix (Continuous Concept Mixing) integrates discrete token prediction with continuous concept vectors derived from sparse autoencoders (SAEs) bricken2023monosemanticity; cunningham2023sparse. By interleaving these continuous representations into the hidden states during training, CoCoMix improves sample efficiency and enriches the model’s capacity for higher-order abstraction, which is beneficial for tasks such as summarization and logical reasoning tack2025llm. Additionally, latent thought vectors (LTMs) were introduced to probabilistically guide token generation via cross-attention mechanisms kong2025scalable. LTMs are sampled from a latent prior and offer a flexible framework that allows adjustments of both inference steps and latent vector dimensions to optimize performance. This method improves sample efficiency and scalability compared to traditional autoregressive models.
Despite these advances, challenges remain. The reliance on implicit reasoning shortcuts may impede true stepwise reasoning, particularly in complex, multi-step tasks lin2025implicitreasoningtransformersreasoning. Future research may focus on refining these latent space pretraining methods to ensure consistency and accuracy across a broader range of reasoning tasks.
@@ 6.2 Subquadratic Attention
Among subquadratic attention mechanisms, linear sequence modeling techniques sun2025linear, such as linear attention, state space models (SSMs), and linear RNNs, emerge as effective alternatives to traditional self-attention. Sparse attention also presents a viable solution by selectively focusing on a subset of tokens, further improving computational efficiency. In the following sections, we will explore these subquadratic attention mechanisms in detail. Additionally, we will discuss recent linearization methods that transform pre-trained transformer-based model weights into linear recurrent structures, allowing for efficient inference while maintaining the knowledge embedded in the original transformer models.